# Principled Foundations for Preference $**{}^{\mbox{\small*}}$

### v1

@@ 4.4 The DPO menagerie in the context of KLST+S
Let $(\ell_{0},\ell_{1})$ be proper (resp. strictly proper), and $a>0$ and $c\in\mathbb{R}$ any constants. Then the loss (~$(\tilde{\ell}_{0},\tilde{\ell}_{1})\stackrel{{\scriptstyle\mathrm{.}}}{{=}}(a\ell_{0},a\ell_{1}+c)$ is proper (resp. strictly proper) and furthermore, using notations from Theorem 7 we have the following key quantities for loss (~$(\tilde{\ell}_{0},\tilde{\ell}_{1})$: $\tilde{H}^{-1}(z)=H^{-1}((z+c)/a)$ and $\tilde{\phi}^{\star}(z)=a\phi^{\star}((z+c)/a)$.
@@
Suppose we wish to compute some $\tilde{\pi}(y_{i}|x)$ that would satisfy two constraints: (a) it is the best probability approximation to all $\{\pi(y_{i,k}|y_{i,<k},x)\}_{i=1}^{n_{i}}$, $\tilde{\pi}(y_{i}|x)$ and (b) it has the form $\tilde{\pi}(y_{i}|x)=u(\pi(y_{i}|x))$ via (20), for some $u$ to specify that would carry some knowledge about $\{\pi(y_{i,k}|y_{i,<k},x)\}_{i=1}^{n_{i}}$ – the simplest of which being the size, $n_{i}$. Plugging $\tilde{\pi}(y_{i}|x)=u(\pi(y_{i}|x))$ in (15) would then carry this knowledge at zero cost in the DPO pipeline. Formulating (a) is just a matter of reusing (11) by using $\pi(y_{..}|y_{..},x)$ as reference policy $\pi_{\mbox{\tiny{ref}}}$, assuming all rewards equal — and of course after Theorem 5, using a Bregman divergence $D_{\phi,\bm{G}}$ as criterion so we end up looking at the argument probability maximizing (for all $i=1,2,...,n$)
@@
        $n=2$ $n>2$
      log     $-\log(p_{i})$     $1/(1+\exp(-z))$     $\log(1+\exp(-z))$     (Rafailov et al., 2023)
    binary entropy     $-\log(p_{i})+\sum_{j\neq i}-\log\left(1-p_{j}\right)$     $1/(1+\exp(-z/2))$     $2\log(1+\exp(-z/2))$     (Hong et al., 2024)
@@ 4.5 Cautionary places and workarounds in the context of KLST+S
    if $L$ is proper, it can be represented as $\bm{\ell}(\bm{q})=U(\bm{q})\cdot\bm{1}+[u_{i}(\bm{q})]_{i}$, such that its ABF representation $D_{\phi}(\bm{p}\|\bm{q})$ satisfies $\phi=-U(\bm{p})+\bm{p}^{\top}(-\bm{u}(\bm{p}))$, $\varphi^{\star}(-\bm{u}(\bm{q}))=U(\bm{q})$ and $-\bm{u}\in\partial\phi$.
@@ VII.1 Proof of Theorem 5
Williamson et al. (2016, Proposition 7) shows the standard connection, but with our formalism, we need to also show that $-\bm{\ell}\in\partial-L$, which is simple but yet needs to be shown for completeness. Properness condition imposes $\bm{p}^{\top}\bm{\ell}(\bm{p})\leq\bm{p}^{\top}\bm{\ell}(\bm{q})$ for all $\bm{p},\bm{q}\in\Delta_{n}$. Put all on the right hand side and add there $-L(\bm{q},\bm{q})+L(\bm{q},\bm{q})$. Expand with the definition of $L$ and it yields:
@@
which implies indeed $\bm{\ell}\in\partial-L$. The convexity of $-L(\bm{q},\bm{q})$ was originally shown in Gneiting and Raftery (2007).
@@ VII.2 Proof of Theorem 6
11 11 $=$ $\alpha^{2}(p({\color[rgb]{1,0,0}y_{1}}\succ{\color[rgb]{1.0,0.88,0.21}y_{3}}|x)+p({\color[rgb]{1.0,0.88,0.21}y_{3}}\succ{\color[rgb]{1,0,0}y_{1}}|x))+\alpha(1-\alpha)(p({\color[rgb]{1,0,0}y_{1}}\succ{\color[rgb]{0,0.8046875,0}y_{4}}|x)+p({\color[rgb]{0,0.8046875,0}y_{4}}\succ{\color[rgb]{1,0,0}y_{1}}|x))$
@@
The crux of our proof consists in bypassing the use of the solvability axiom of the original proof (Krantz et al., 1989, Chapter 17, Theorem 2) – which is hard to motivate in the context of preference optimization – by creating a structure on top of alternatives / actions which can be (i) motivated in the context of preference optimization and (ii) compatible with the sampling process of preferences. We have resorted to simple – binary – lotteries, which have been motivated in a broad social choice context for decades (Machina, 1985). Other structures could be eligible. The major "modelling" price to pay is the definition of a local choice structure (LCS) that relies on properties that need to hold for all non-trivial $\alpha$s. This is a stark contrast with monotonicity, which just needs to hold for one. We could have alleviated considering all of $(0,1)$ for the LCS by considering two open subsets of $[0,1]$, one close to 0 and one close to 1. The formulation of a LCS we adopted is more restrictive but more readable, so we kept it.
@@ VII.3 Proof of Theorem 7
Now, if the loss is symmetric, then it has $\ell_{0}(p)=\ell_{1}(1-p)$, inducing the symmetry $H(1-p)=-H(p)$, yielding
@@
and so $\ell\circ H^{-1}(x)=\phi^{\star}(-x)$ (with the convention =.$\ell\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\ell_{1}$), as claimed.
@@ VII.4 Proof of Theorem 11
as claimed. To prove the Theorem, we just have to remark that =.$\ell\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\psi\circ F^{-1}$ is a function $[0,1]\rightarrow\mathbb{R}$ and strictly increasing, being the composition of two strictly increasing functions, which concludes using Lemma Y.
@@ VII.5 Proof of Theorem 15
We want $-\phi(p)=p\cdot\ell_{1}(p)+(1-p)\cdot\ell_{0}(p)$ for some strictly proper loss $(\ell_{0},\ell_{1})$, that we assume wlog continuous on $(0,1)$ (because $\phi$ is necessarily continuous on $(0,1)$, being convex). We get the necessary relationship for $p<1$:
@@ VII.7 Proof of Theorem 21
so $\phi$ has to be convex on $\Delta_{n}$. The target $\bm{p}$ has three or more non-zero probabilities in indexes $\mathbb{I}$. In this case, if we restrict $\bm{q}$ to have non-zero coordinates in the same indexes (allowed by the properness assumption), then any two coordinates of $\bm{q}$ on these indexes are allowed to take values in a non-empty 2-dimensional open subset of a vector space; this, along with convexity and (35), impose the subdifferential property $-\bm{\ell}\in\partial\phi$ (without loss of generality, we suppose vectors restricted to indexes whose coordinates in $\bm{p}$ are non zero) which, since $\phi$ is separable and of the form $\phi(\bm{p})=-\bm{p}^{\top}\bm{\ell}(\bm{p})$, provides at least 3 properties:
@@
Part (I) has crossed almost seven decades under different forms with more restrictive assumption that would have limited its scope in our context: (McCarthy, 1956) mentions an unpublished result imposing all $\ell_{i}$s to be the same; the proof of (Savage, 1971, Section 9.4) imposes strict properness and differentiability on $\bm{\ell}$, and show the result for vectors $\bm{p}$ partially uniform; (Bernardo, 1979, Theorem 2) imposes smoothness, full support and all $\ell_{i}$s to be the same; Dawid (2007) imposes differentiability and all $\ell_{i}$s to be the same. While it has the benefit of covering all proper losses, our generalization is formulated to insist on the fast that there is no "lucky" target $\bm{p}$ for which the loss could be different from the KL divergence in Part (II) [^6] Clearly, $\mathrm{Card}(\mathbb{I})>2$ in RL, even for the simplest of toy problems (Xu et al., 2024b)..
@@
We did not put any constraint on the decomposition of loss coordinates but (40) gives us a very clear one: it must be done in such a way that (i) $-U(\bm{q})+\bm{q}^{\top}(-\bm{u}(\bm{q}))$ is convex (see Part I) and (ii)
@@ VII.8 Proof of Lemma 22
where we have used $n$ times the properness of $(\ell_{0},\ell_{1})$ in the inequality (strict properness would involve a strict inequality). The last equality demonstrates the simplicity of the final loss in (17) ($L(\bm{p},\bm{p})=\sum_{i}p_{i}\ell_{0}(p_{i})+(1-p_{i})\ell_{1}\left(1-p_{i}\right)$), as exemplified by the binary entropy in ORPO (Hong et al., 2024).

### v2

@@ 4.4 The DPO menagerie in the context of KLST+S
Let $(\ell_{0},\ell_{1})$ be proper (resp. strictly proper), and $a>0$ and $c\in\mathbb{R}$ any constants. Then the loss $(\tilde{\ell}_{0},\tilde{\ell}_{1})\stackrel{{\scriptstyle\mathrm{.}}}{{=}}(a\ell_{0},a\ell_{1}+c)$ is proper (resp. strictly proper) and furthermore, using notations from Theorem 7 we have the following key quantities for loss $(\tilde{\ell}_{0},\tilde{\ell}_{1})$: $\tilde{H}^{-1}(z)=H^{-1}((z+c)/a)$ and $\tilde{\phi}^{\star}(z)=a\phi^{\star}((z+c)/a)$.
@@
Suppose we wish to compute some $\tilde{\pi}(y_{i}|x)$ that would satisfy two constraints: (a) it is the best probability approximation to all $\{\pi(y_{i,k}|y_{i,<k},x)\}_{i=1}^{n_{i}}$, $\tilde{\pi}(y_{i}|x)$ and (b) it has the form $\tilde{\pi}(y_{i}|x)=u(\pi(y_{i}|x))$ via (21), for some $u$ to specify that would carry some knowledge about $\{\pi(y_{i,k}|y_{i,<k},x)\}_{i=1}^{n_{i}}$ – the simplest of which being the size, $n_{i}$. Plugging $\tilde{\pi}(y_{i}|x)=u(\pi(y_{i}|x))$ in (16) would then carry this knowledge at zero cost in the DPO pipeline. Formulating (a) is just a matter of reusing (11) by using $\pi(y_{..}|y_{..},x)$ as reference policy $\pi_{\mbox{\tiny{ref}}}$, assuming all rewards equal — and of course after Theorem 5, using a Bregman divergence $D_{\phi,\bm{G}}$ as criterion so we end up looking at the argument probability maximizing (for all i=1,2,,ni=1,2,...,n)
@@
        $n=2$ ? $n>2$ ?
      log     $\log(p_{i})$     $1/(1+\exp(-z))$     $\log(1+\exp(-z))$     (Rafailov et al., 2023)
    binary entropy     $\log(p_{i})+\sum_{j\neq i}-\log\left(1-p_{j}\right)$     $1/(1+\exp(-z/2))$     $2\log(1+\exp(-z/2))$     (Hong et al., 2024)
@@ 4.5 Cautionary places and workarounds in the context of KLST+S
    if $L$ is proper, it can be represented as $\bm{\ell}(\bm{q})=U(\bm{q})\cdot\bm{1}+[u_{i}(\bm{q})]_{i}$, such that its ABF representation $D_{\phi}(\bm{p}\|\bm{q})$ satisfies $\phi=-U(\bm{p})+\bm{p}^{\top}(-\bm{u}(\bm{p}))$, $\varphi^{\star}(-\bm{u}(\bm{q}))=U(\bm{q})$ and $\bm{u}\in\partial\phi$.
@@ 5 Limitations, conclusion and open problems
Acknowledgments
The authors warmly thank Robert C. Williamson for many comments and discussions about this material.
@@ VII.1 Proof of Theorem 5
Williamson et al. (2016, Proposition 7) shows the standard connection, but with our formalism, we need to also show that $\bm{\ell}\in\partial-L$, which is simple but yet needs to be shown for completeness. Properness condition imposes $\bm{p}^{\top}\bm{\ell}(\bm{p})\leq\bm{p}^{\top}\bm{\ell}(\bm{q})$ for all $\bm{p},\bm{q}\in\Delta_{n}$. Put all on the right hand side and add there $L(\bm{q},\bm{q})+L(\bm{q},\bm{q})$. Expand with the definition of $L$ and it yields:
@@
which implies indeed $\bm{\ell}\in\partial-L$. The convexity of $L(\bm{q},\bm{q})$ was originally shown in Gneiting and Raftery (2007).
@@ VII.2 Proof of Theorem 6
$1$ $=$ $\alpha^{2}(p({\color[rgb]{1,0,0}y_{1}}\succ{\color[rgb]{1.0,0.88,0.21}y_{3}}|x)+p({\color[rgb]{1.0,0.88,0.21}y_{3}}\succ{\color[rgb]{1,0,0}y_{1}}|x))+\alpha(1-\alpha)(p({\color[rgb]{1,0,0}y_{1}}\succ{\color[rgb]{0,0.8046875,0}y_{4}}|x)+p({\color[rgb]{0,0.8046875,0}y_{4}}\succ{\color[rgb]{1,0,0}y_{1}}|x))$
@@
[…]
@@ VII.3 Proof of Theorem 7
Now, if the loss is symmetric, then it has $\ell_{0}(p)=\ell_{1}(1-p)$, inducing the symmetry H(1p)=H(p)H(1-p)=-H(p), yielding
@@
and so $\ell\circ H^{-1}(x)=\phi^{\star}(-x)$ (with the convention $\ell\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\ell_{1}$), as claimed.
@@ VII.4 Proof of Theorem 11
as claimed. To prove the Theorem, we just have to remark that $\ell\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\psi\circ F^{-1}$ is a function $0,1]\rightarrow\mathbb{R}$ and strictly increasing, being the composition of two strictly increasing functions, which concludes using Lemma Y.
@@ VII.5 Proof of Theorem 15
We want $\phi(p)=p\cdot\ell_{1}(p)+(1-p)\cdot\ell_{0}(p)$ for some strictly proper loss $(\ell_{0},\ell_{1})$, that we assume wlog continuous on $(0,1)$ (because $\phi$ is necessarily continuous on $(0,1)$, being convex). We get the necessary relationship for $p<1$:
@@ VII.7 Proof of Theorem 21
so $\phi$ has to be convex on $\Delta_{n}$. The target $\bm{p}$ has three or more non-zero probabilities in indexes $\mathbb{I}$. In this case, if we restrict $\bm{q}$ to have non-zero coordinates in the same indexes (allowed by the properness assumption), then any two coordinates of $\bm{q}$ on these indexes are allowed to take values in a non-empty 2-dimensional open subset of a vector space; this, along with convexity and (36), impose the subdifferential property $\bm{\ell}\in\partial\phi$ (without loss of generality, we suppose vectors restricted to indexes whose coordinates in $\bm{p}$ are non zero) which, since $\phi$ is separable and of the form $\phi(\bm{p})=-\bm{p}^{\top}\bm{\ell}(\bm{p})$, provides at least 3 properties:
@@
[…]
@@
We did not put any constraint on the decomposition of loss coordinates but (41) gives us a very clear one: it must be done in such a way that (i) $U(\bm{q})+\bm{q}^{\top}(-\bm{u}(\bm{q}))$ is convex (see Part I) and (ii)
@@ VII.8 Proof of Lemma 22
[…]